Single plantwide factory overhead rate

Bach Instruments Inc. makes three musical instruments: flutes, clarinets, and oboes. The budgeted factory overhead cost is $94,200. Overhead is allocated to the three products on the basis of direct labor hours. The products have the following budgeted production volume and direct labor hours per unit:

Budgeted Production Volume Direct Labor Hours Per Unit
Flutes 1,900 units 0.4
Clarinets 700 1.4
Oboes 1,400 1.0

If required, round all per unit answers to the nearest cent.

a. Determine the single plantwide overhead rate.
$ per direct labor hour

b. Use the overhead rate in (a) to determine the amount of total and per-unit overhead allocated to each of the three products.

Total
Factory Overhead Cost
Per Unit
Factory Overhead Cost
Flutes $ $
Clarinets
Oboes
Total $
Answer #1

А B с D Total direct labor hours 760 980 1400 3140 1 a) Budgeted Direct labor hours 2 production per unit 3 Flutes 1,900 0.4

с D w N Direct labor hours per unit 0.4 Total direct labor hours =+B3*C3 =+B4*C4 =+B5*C5 =+D3+04+05 1.4 1 A B 1 a) Budgeted p